# Universal Physics Transformers: A Framework For Efficiently Scaling Neural Operators

### v2

@@ A.1 Navier-Stokes equations
$\bm{u}\cdot\nabla\bm{u}=\left(\begin{array}[]{cc}u_{1}&u_{2}\\ \end{array}\right)\cdot\left(\begin{array}[]{cc}\dfrac{\partial u_{1}}{\partial x}&\dfrac{\partial u_{2}}{\partial x}\\[8.61108pt] \dfrac{\partial u_{1}}{\partial y}&\dfrac{\partial u_{2}}{\partial y}\end{array}\right)$ (8)
@@
$\mu\nabla^{2}\bm{u}=\mu\left(\begin{array}[]{cc}\dfrac{\partial}{\partial x}&\dfrac{\partial}{\partial y}\\ \end{array}\right)\cdot\left(\begin{array}[]{c}\dfrac{\partial}{\partial x}\\[8.61108pt] \dfrac{\partial}{\partial y}\\ \end{array}\right)\left(\begin{array}[]{c}u_{1}\\ u_{2}\\ \end{array}\right)=\mu\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)\left(\begin{array}[]{c}u_{1}\\ u_{2}\\ \end{array}\right)$ (16)
@@
$\left(\begin{array}[]{cc}\dfrac{\partial}{\partial x}&\dfrac{\partial}{\partial y}\\ \end{array}\right)\cdot\left(\begin{array}[]{c}u_{1}\\ u_{2}\\ \end{array}\right)=\frac{\partial u_{1}}{\partial x}+\frac{\partial u_{2}}{\partial y}=0$ (20)
@@
A.2 Turbulence
Turbulence (Pope, 2001; Mathieu & Scott, 2000) is one of the key aspects of CFD and refers to chaotic and irregular motion of fluid flows, such as air or water. It is characterized by unpredictable changes in velocity, pressure, and density within the fluid. Turbulent flow is distinguished from laminar flow, which is smooth and orderly. There are several factors that can contribute to the onset of turbulence, including high flow velocities, irregularities in the shape of surfaces over which the fluid flows, and changes in the fluid’s viscosity. Turbulence plays a significant role in many natural phenomena and engineering applications, influencing processes such as mixing, heat transfer, and the dispersion of particles in fluids. Understanding and predicting turbulence is crucial in fields like fluid dynamics, aerodynamics, and meteorology. Turbulent flows occur at high Reynolds numbers
@@ A.2 Turbulence
where $\rho$ is the density of the fluid, $\nu$ is the velocity of the fluid, $L$ is the linear dimension, e.g., the width of a pipe, and $\mu$ is the viscosity parameter. Turbulent flows and are dominated by inertial forces, which tend to produce chaotic eddies, vortices and other flow instabilities.
Turbulence in 3D vs turbulence in 2D. For high Reynolds numbers, i.e., in the turbulent regime, solutions to incompressible Navier-Stokes equations differ in three dimensions compared to two dimensions by a phenomenon called energy cascade. Energy cascade orchestrates the transfer of kinetic energy from large-scale vortices to progressively smaller scales until it is eventually converted into thermal energy. In three dimensions, this energy cascade is responsible for to the continuous formation of vortex structures at various scales, and thus for the emergence of high-frequency features. In contrast to three dimensions, the energy cascade is inverted in two dimensions, i.e., energy is transported from smaller to larger scales, resulting in more homogeneous, long-lived structures. Mathematically the difference between three dimensional and two dimensional turbulence modeling can be best seen be rewriting the Navier-Stokes equations in vorticity formulation, where the vorticity ${\bm{\omega}}$ is the curl of the flow field, i.e., ${\bm{\omega}}=\nabla\times\bm{u}$, with $\times$ here denoting the cross product. This derivation can be found in many standard texts on Navier-Stokes equation , e.g. Tsai (2018, Section 1.4.), and for the reader’s convenience we briefly repeat the derivation here. We start using Eq. (4), setting $\bm{f}=0$ for simplicity:
@@
where $\frac{D{\bm{\omega}}}{Dt}=\frac{\partial{\bm{\omega}}}{\partial t}+\left(\bm{u}\cdot\nabla\right){\bm{\omega}}$.
For two dimensional flows
$\bm{u}=\left(\begin{array}[]{c}u_{1}\\ u_{2}\\ 0\\ \end{array}\right)\ \text{and}\ \ {\bm{\omega}}=\nabla\times\bm{u}=\left(\begin{array}[]{c}0\\ 0\\ \frac{\partial u_{2}}{\partial x}-\frac{\partial u_{1}}{\partial y}\\ \end{array}\right)$ (35)
and thus $\frac{\partial{\bm{\omega}}}{\partial t}=0$ in two dimensions. Consequently, the length of a vortex cannot change in two dimensions, resulting in homogeneous, long-lived structures. Conversely, in three dimensions,
$\frac{D\omega_{i}}{Dt}=\underbrace{\omega_{j}\frac{\partial u_{i}}{\partial x_{j}}}_{\text{vortex turning and stretching}}+\underbrace{\mu\frac{\partial^{2}\omega_{i}}{\partial x_{j}\partial x_{j}}}_{\text{diffusion}}\ .$ (36)
For example, picking an arbitrary component:
@@
There’s been several approaches to model turbulence with the help of Deep Learning. Predominantly models for two dimensional scenarios have been suggested (Pfaff et al., 2020; Li et al., 2020a; Brandstetter et al., 2022c; Kochkov et al., 2021; Kohl et al., 2023). Due to the higher complexity (as discussed above) and memory- and compute costs comparably less work was done in the 3D case (Stachenfeld et al., 2022; Lienen et al., 2023).
A.3 Numerical modeling based on Eulerian discretization schemes
Various approaches are available for numerically addressing turbulence, with different schemes suited to different levels of computational intensity. Among these, Direct Numerical Simulation (DNS) (Moin & Mahesh, 1998) stands out as the most resource-intensive method, involving the direct solution of the unsteady Navier-Stokes equations.
DNS is renowned for its capability to resolve even the minutest eddies and time scales present in turbulent flows. While DNS does not necessitate additional closure equations, a significant drawback is its high computational demand. Achieving accurate solutions requires the utilization of very fine grids and extremely small time steps. This is based on Kolmoqorov (1941) and Kolmogorov (1962) which give the minimum spacial and temporal scales that need to be resolved for accurately simulating turbulence. The discretization-scales for both time and space needed for 3D-problems become very small and simulations computationally extremely expensive.
This computational intensity of DNS resulted in the development of turbulence modeling, with Large Eddy Simulations (LES) (Smagorinsky, 1963) and Reynolds-averaged Navier-Stokes equations (RANS) (Reynolds, 1895) being two prominent examples. LES aim to reduce computational cost by neglecting the computationally expensive smallest length scales in turbulent flows. This is achieved through a low-pass filtering of the Navier–Stokes equations, effectively removing fine-scale details via time- and spatial-averaging. In contrast, the Reynolds-averaged Navier–Stokes equations (RANS equations) are based on time-averaging. The foundational concept is Reynolds decomposition, attributed to Osborne Reynolds, where an instantaneous quantity is decomposed into its time-averaged and fluctuating components.
@@ A.3 Numerical modeling based on Eulerian discretization schemes
A.3.1 Reynolds-averaged Navier-Stokes equations
@@
Using the the mass conserving equation (A.3.1) the momentum equation (39) can be rewritten as
@@
where $-\overline{u^{\prime}_{i}u^{\prime}_{j}}$ is called the Reynolds stress which needs further modeling to solve the above equations. More specifically, this is referred to as the Closure Problem which led to many turbulence models such as $k\text{-}\epsilon$ (Hanjalic & Brian, 1972) or $k\text{-}\omega$ (Wilcox, 2008). Analogously the 3D RANS equations can be derived. This turbulence model consists of simplified equations that predict the statistical evolution of turbulent flows. Due to the Reynolds stress there still remain velocity fluctuations in the RANS equations. To get equations that contain only time-averaged quantities the RANS equations need to be closed by modeling the Reynolds stress as a function of the mean flow such that any reference to the fluctuating parts is removed. The first such approach led to the eddy viscosity model (Boussinesq, 1877) for 3-d incompressible Navier Stokes:
@@
A.3.2 SIMPLE and PISO algorithm
Next, we introduce two popular algorithms that try to solve Eq. (4) numerically, where the main idea is to couple pressure and velocity computations. Specifically, we will discuss the SIMPLE (Semi-Implicit Method for Pressure-Linked Equations) (Patankar & Spalding, 1972) and PISO (Pressure Implicit with Splitting of Operators) (Issa, 1986) algorithm, which are implemented in OpenFOAM as well. The essence of these algorithms are the following four main steps:
@@ A.3.2 SIMPLE and PISO algorithm
A.4 Lagrangian discretization schemes
In contrast to such grid- and mesh-based representations, in Lagrangian schemes, the discretization is carried out using finitely many material points, often referred to as particles, which move with the local deformation of the continuum. Roughly speaking, there are three families of Lagrangian schemes: discrete element methods (Cundall & Strack, 1979), material point methods (Sulsky et al., 1994; Brackbill & Ruppel, 1986), and smoothed particle hydrodynamics (SPH) (Gingold & Monaghan, 1977; Lucy, 1977; Monaghan, 1992, 2005).
A.4.1 Smoothed particle hydrodynamics
@@
in terms of these kernel interpolations leads to a system of ordinary differential equations (ODEs) for the particle accelerations (Morris et al., 1997) where the respective velocities and positions can be computed by integration. One of the advantages of SPH compared to Eulerian discretization techniques is that SPH can handle large topological changes as no connectivity constraints between particles are required, and advection is treated exactly (Toshev et al., 2023a; Monaghan & Gingold, 1983). The lack of a mesh significantly simplifies the model implementation and opens up more possibilities for parallelization compared to Eulerian schemes (Harada et al., 2007; Crespo et al., 2011). However, for accurately resolving particle dynamics on small scales a large number of particles is needed to achieve a resolution comparable to grid-based methods specifically when the metric of interest is not directly related to density which makes SPH more expensive (Colagrossi & Maurizio, 2003). Also, setting boundary conditions such as inlets, outlets, and walls is more difficult than for grid-based methods.
A seemingly great fit for particle-based dynamics are graph neural networks (GNNs) (Scarselli et al., 2008; Kipf & Welling, 2017) with graph-based latent space representations. In many cases, predicted accelerations at the nodes are numerically integrated to model the time evolution of the many-particle systems (Sanchez-Gonzalez et al., 2020; Mayr et al., 2023; Toshev et al., 2023a, 2024).
Appendix B Experimental details
B.1 General
All experiments are conducted on A100 GPUs. All experiments linearly scale the learning rate with the batchsize, exclude normalization and bias parameters from the weight decay, follow a linear warmup $\rightarrow$ cosine decay learning rate schedule (Loshchilov & Hutter, 2017) and use the AdamW (Kingma & Ba, 2015; Loshchilov & Hutter, 2019) optimizer. Transformer blocks follow a standard pre-norm architecture as used in ViT (Dosovitskiy et al., 2021).
@@ B.1 General
B.2 Feature modulation
We apply feature modulation to condition models to time/velocity. We use different forms of feature modulation in accordance with existing works depending on the model type. FiLM (Perez et al., 2018) is used for U-Net, the “Spatial-Spectral” conditioning of Gupta & Brandstetter (2022) is used for FNO based models and DiT (Peebles & Xie, 2023) is used for transformer blocks. For perceiver blocks, we extend the modulation of DiT to seperately modulate queries from keys/values.
B.3 ShapeNet-Car

### v3

@@ B.1 Navier-Stokes equations
$\bm{u}\cdot\nabla\bm{u}=\left(\begin{array}[]{cc}u_{1}&u_{2}\\ \end{array}\right)\cdot\left(\begin{array}[]{cc}\dfrac{\partial u_{1}}{\partial x_{1}}&\dfrac{\partial u_{2}}{\partial x_{1}}\\[8.61108pt] \dfrac{\partial u_{1}}{\partial x_{2}}&\dfrac{\partial u_{2}}{\partial x_{2}}\end{array}\right)$ (8)
@@
$\mu\nabla^{2}\bm{u}=\mu\left(\begin{array}[]{cc}\dfrac{\partial}{\partial x_{1}}&\dfrac{\partial}{\partial x_{2}}\\ \end{array}\right)\cdot\left(\begin{array}[]{c}\dfrac{\partial}{\partial x_{1}}\\[8.61108pt] \dfrac{\partial}{\partial x_{2}}\\ \end{array}\right)\left(\begin{array}[]{c}u_{1}\\ u_{2}\\ \end{array}\right)=\mu\left(\frac{\partial^{2}}{\partial x_{1}^{2}}+\frac{\partial^{2}}{\partial x_{2}^{2}}\right)\left(\begin{array}[]{c}u_{1}\\ u_{2}\\ \end{array}\right)$ (16)
@@
$\left(\begin{array}[]{cc}\dfrac{\partial}{\partial x_{1}}&\dfrac{\partial}{\partial x_{2}}\\ \end{array}\right)\cdot\left(\begin{array}[]{c}u_{1}\\ u_{2}\\ \end{array}\right)=\frac{\partial u_{1}}{\partial x_{1}}+\frac{\partial u_{2}}{\partial x_{2}}=0$ (20)
@@
B.2 Turbulence
Turbulence [82, 64] is one of the key aspects of CFD and refers to chaotic and irregular motion of fluid flows, such as air or water. It is characterized by unpredictable changes in velocity, pressure, and density within the fluid. Turbulent flow is distinguished from laminar flow, which is smooth and orderly. There are several factors that can contribute to the onset of turbulence, including high flow velocities, irregularities in the shape of surfaces over which the fluid flows, and changes in the fluid’s viscosity. Turbulence plays a significant role in many natural phenomena and engineering applications, influencing processes such as mixing, heat transfer, and the dispersion of particles in fluids. Understanding and predicting turbulence is crucial in fields like fluid dynamics, aerodynamics, and meteorology. Turbulent flows occur at high Reynolds numbers
@@ B.2 Turbulence
where $\rho$ is the density of the fluid, $\nu$ is the velocity of the fluid, $L$ is the linear dimension, e.g., the width of a pipe, and $\mu$ is the viscosity parameter. Turbulent flows are dominated by inertial forces, which tend to produce chaotic eddies, vortices and other flow instabilities.
Turbulence in 3D vs turbulence in 2D
For high Reynolds numbers, i.e., in the turbulent regime, solutions to incompressible Navier-Stokes equations differ in three dimensions compared to two dimensions by a phenomenon called energy cascade. Energy cascade orchestrates the transfer of kinetic energy from large-scale vortices to progressively smaller scales until it is eventually converted into thermal energy. In three dimensions, this energy cascade is responsible for the continuous formation of vortex structures at various scales, and thus for the emergence of high-frequency features. In contrast to three dimensions, the energy cascade is inverted in two dimensions, i.e., energy is transported from smaller to larger scales, resulting in more homogeneous, long-lived structures. Mathematically the difference between three dimensional and two dimensional turbulence modeling can be best seen by rewriting the Navier-Stokes equations in vorticity formulation, where the vorticity ${\bm{\omega}}$ is the curl of the flow field, i.e., ${\bm{\omega}}=\nabla\times\bm{u}$, with $\times$ here denoting the cross product. This derivation can be found in many standard texts on Navier-Stokes equation , e.g. Tsai [104, Section 1.4.], and for the reader’s convenience we briefly repeat the derivation here. We start using Eq. (4), setting $\bm{f}=0$ for simplicity:
@@
with the material derivative $\frac{D{\bm{\omega}}}{Dt}=\frac{\partial{\bm{\omega}}}{\partial t}+\left(\bm{u}\cdot\nabla\right){\bm{\omega}}$.
For three dimensional flows the material derivative of each component $\omega_{i}$ can be written as
$\frac{D\omega_{i}}{Dt}=\sum_{j=1}^{3}\left(\underbrace{\omega_{j}\frac{\partial u_{i}}{\partial x_{j}}}_{\text{vortex turning and stretching}}+\underbrace{\mu\frac{\partial^{2}\omega_{i}}{\partial x_{j}\partial x_{j}}}_{\text{diffusion}}\right)\ ,$ (29)
for example picking $i=2$:
@@
For two-dimensional flows, as represented by
$\bm{u}=\left(\begin{array}[]{c}u_{1}\\ u_{2}\\ 0\\ \end{array}\right)\ \text{and}\ \ {\bm{\omega}}=\nabla\times\bm{u}=\left(\begin{array}[]{c}0\\ 0\\ \frac{\partial u_{2}}{\partial x_{1}}-\frac{\partial u_{1}}{\partial x_{2}}\\ \end{array}\right)\ ,$ (37)
the terms for vortex turning and vortex stretching vanish, as is evident in the material derivative of $\omega_{3}$, given by
$\frac{D\omega_{3}}{Dt}=\underbrace{\omega_{1}\frac{\partial u_{3}}{\partial x_{1}}}_{=0}+\underbrace{\omega_{2}\frac{\partial u_{3}}{\partial x_{2}}}_{=0}+\underbrace{\omega_{3}\frac{\partial u_{3}}{\partial x_{3}}}_{=0}+\underbrace{\mu\left(\frac{\partial^{2}\omega_{3}}{\partial x_{1}\partial x_{1}}+\ \frac{\partial^{2}\omega_{3}}{\partial x_{2}\partial x_{2}}\right)}_{\text{diffusion}}+\underbrace{\mu\frac{\partial^{2}\omega_{3}}{\partial x_{3}\partial x_{3}}}_{=0}.$ (38)
Consequently, the length and the angle of a vortex do not change in two dimensions, resulting in homogeneous, long-lived structures.
There’s been several approaches to model turbulence with the help of Deep Learning. Predominantly models for two dimensional scenarios have been suggested [81, 51, 15, 43, 44]. Due to the higher complexity (as discussed above) and memory- and compute costs comparably less work was done in the 3D case [92, 56].
B.3 Numerical modeling based on Eulerian discretization schemes
Various approaches are available for numerically addressing turbulence, with different schemes suited to different levels of computational intensity. Among these, Direct Numerical Simulation (DNS) [68] stands out as the most resource-intensive method, involving the direct solution of the unsteady Navier-Stokes equations.
DNS is renowned for its capability to resolve even the minutest eddies and time scales present in turbulent flows. While DNS does not necessitate additional closure equations, a significant drawback is its high computational demand. Achieving accurate solutions requires the utilization of very fine grids and extremely small time steps. This is based on Kolmoqorov [46] and Kolmogorov [45] which give the minimum spacial and temporal scales that need to be resolved for accurately simulating turbulence. The discretization-scales for both time and space needed for 3D-problems become very small and simulations computationally extremely expensive.
This computational intensity of DNS resulted in the development of turbulence modeling, with Large Eddy Simulations (LES) [91] and Reynolds-averaged Navier-Stokes equations (RANS) [84] being two prominent examples. LES aim to reduce computational cost by neglecting the computationally expensive smallest length scales in turbulent flows. This is achieved through a low-pass filtering of the Navier–Stokes equations, effectively removing fine-scale details via time- and spatial-averaging. In contrast, the Reynolds-averaged Navier–Stokes equations (RANS equations) are based on time-averaging. The foundational concept is Reynolds decomposition, attributed to Osborne Reynolds, where an instantaneous quantity is decomposed into its time-averaged and fluctuating components.
@@ B.3 Numerical modeling based on Eulerian discretization schemes
B.3.1 Reynolds-averaged Navier-Stokes equations
@@
Using the the mass conserving equation (B.3.1) the momentum equation (40) can be rewritten as
@@
where $-\overline{u^{\prime}_{i}u^{\prime}_{j}}$ is called the Reynolds stress which needs further modeling to solve the above equations. More specifically, this is referred to as the Closure Problem which led to many turbulence models such as $k\text{-}\epsilon$ [32] or $k\text{-}\omega$ [112]. Analogously the 3D RANS equations can be derived. This turbulence model consists of simplified equations that predict the statistical evolution of turbulent flows. Due to the Reynolds stress there still remain velocity fluctuations in the RANS equations. To get equations that contain only time-averaged quantities the RANS equations need to be closed by modeling the Reynolds stress as a function of the mean flow such that any reference to the fluctuating parts is removed. The first such approach led to the eddy viscosity model [11] for 3-d incompressible Navier Stokes:
@@
B.3.2 SIMPLE and PISO algorithm
Next, we introduce two popular algorithms that try to solve Eq. (4) numerically, where the main idea is to couple pressure and velocity computations. Specifically, we will discuss the SIMPLE (Semi-Implicit Method for Pressure-Linked Equations) [77] and PISO (Pressure Implicit with Splitting of Operators) [38] algorithm, which are implemented in OpenFOAM as well. The essence of these algorithms are the following four main steps:
@@ B.3.2 SIMPLE and PISO algorithm
B.4 Lagrangian discretization schemes
In contrast to such grid- and mesh-based representations, in Lagrangian schemes, the discretization is carried out using finitely many material points, often referred to as particles, which move with the local deformation of the continuum. Roughly speaking, there are three families of Lagrangian schemes: discrete element methods [24], material point methods [94, 12], and smoothed particle hydrodynamics (SPH) [29, 62, 69, 70].
B.4.1 Smoothed particle hydrodynamics
@@
in terms of these kernel interpolations leads to a system of ordinary differential equations (ODEs) for the particle accelerations [72] where the respective velocities and positions can be computed by integration. One of the advantages of SPH compared to Eulerian discretization techniques is that SPH can handle large topological changes as no connectivity constraints between particles are required, and advection is treated exactly [100, 71]. The lack of a mesh significantly simplifies the model implementation and opens up more possibilities for parallelization compared to Eulerian schemes [35, 23]. However, for accurately resolving particle dynamics on small scales a large number of particles is needed to achieve a resolution comparable to grid-based methods specifically when the metric of interest is not directly related to density which makes SPH more expensive [22]. Also, setting boundary conditions such as inlets, outlets, and walls is more difficult than for grid-based methods.
A seemingly great fit for particle-based dynamics are graph neural networks (GNNs) [88, 42] with graph-based latent space representations. In many cases, predicted accelerations at the nodes are numerically integrated to model the time evolution of the many-particle systems [87, 65, 100, 102].
Appendix C Justification that UPTs are universal neural operators
We provide a brief sketch of how universality can be established for our architecture. For transformer-based neural operators, universal approximation has been recently demonstrated in [17], Section 5. The arguments in this work are heavily based on [49], which establishes that nonlinearity and nonlocality are crucial for universality. By demonstrating that the attention mechanism can, under appropriate weight choices, function as an averaging operator, the results from [49] are directly applicable. For detailed proof, refer to Theorem 22 in [17]. Our algorithm fits within this framework as well: we employ nonlinear, attention-based encoders and decoders (as allowed by the results of [49], Section 2) and utilize attention layers in the latent space.
Appendix D Experimental details and extended results
D.1 General
All experiments are conducted mostly on A100 GPUs. For large-scale experiments we use two research clusters equipped with either 8xA100-40GB nodes or 4xA100-64GB nodes. For smaller experiments and evaluations, we use a mix of internal servers equipped with varying numbers of A100-40GB or A100-80GB cards.
@@ D.1 General
All experiments linearly scale the learning rate with the batchsize, exclude normalization and bias parameters from the weight decay, follow a linear warmup $\rightarrow$ cosine decay learning rate schedule [58] and use the AdamW [41, 59] optimizer. Transformer blocks follow a standard pre-norm architecture as used in ViT [26].
D.2 Feature modulation
We apply feature modulation to condition models to time/velocity. We use different forms of feature modulation in accordance with existing works depending on the model type. FiLM [80] is used for U-Net, the “Spatial-Spectral” conditioning of [31] is used for FNO based models and DiT [78] is used for transformer blocks. For perceiver blocks, we extend the modulation of DiT to seperately modulate queries from keys/values.
D.3 Experiments on regular grid datasets
We use two datasets from [31], the “Navier-Stokes 2D” and “ShallowWater-2D” to compare against baseline transformer architectures. Baseline results, data preprocessing and benchmark setups are taken from DPOT [34] and CViT [108].
As these datasets contain only regular grid data, we make the following modifications to UPT: (i) we replace the supernode pooling with a patch embedding (ii) we remove the perceiver pooling in the encoder (iii) we decode patchwise. These modifications were done as (i) patch embedding is more efficient than supernode pooling for regular grid data (ii) these datasets are small-scale datasets, so compressing the latent space to make the model more efficient is not necessary (iii) as the decoding positions are always the same, there is no reason to decode position-wise as the missig variability of positions in the dataset makes superresolution impossible and patch-wise decoding is more efficient than position-wise decoding. This setup is conceptually similar to CViT [108].
We use a peak learning rate of 1e-4 with 10% warmup and a cosine decay schedule afterwards, patchsize 8, batchsize 256 and train for 1000 epochs. We use hidden dimension 96 and 12 blocks for 2M parameter models and hidden dimension 256 with 14 blocks for 13M parameter models to best match model sizes of previous models. Additionally, we mask out between 0% and 5% of the input patches via an attention mask in the transformer and perceiver blocks. Similar performances could be achieved with less epochs, but our flexible masking strategy allows long training without overfitting.
D.3.1 NavierStokes-2D dataset
We run comparisons against different Transformer baselines on a regular gridded Navier-Stokes equations dataset [31] in Table 7(b) (baseline results and evaluation protocol taken from DPOT [34]). UPT outperforms all compared methods, some of which are specifically designed for regularly gridded data.
As baseline transformers often train small models, we compare on a small scale, where UPT significantly outperforms other models. We also compare on larger scales, where UPT again outperforms competitors.
Table 2: Comparison on a regular grid Navier-Stokes dataset [31]. (a) UPT outperforms competitor methods that train only small models by a large margin. (b) UPT also performs well on larger model sizes, outperforming competitors even if they train much larger models or pre-train (PT) on more data followed by fine-tuning (FT) on the Navier-Stokes dataset.
Model # Params Rel. L2 Error
FNO [51] 0.5M 9.12%
FFNO [103] 1.3M 8.39%
GK-T [18] 1.6M 9.52%
GNOT [33] 1.8M 17.20%
OFormer [53] 1.9M 13.50%
UPT-T 1.8M 5.08%
Model # Params Rel. L2 Error
DPOT-Ti [34] 7M 12.50%
DPOT-S [34] 30M 9.91%
DPOT-L (PT) 500M 7.98%
DPOT-L (FT) 500M 2.78%
DPOT-H (PT) 1.03B 3.79%
CViT-S [108] 13M 3.75%
CViT-B [108] 30M 3.18%
UPT-S 13M 3.12%
UPT-B 30M 2.69%
D.3.2 ShallowWater-2D dataset
We run comparisons against UNet, FNO, Dilated ResNet variants on the regular gridded ShallowWater-2D climate modeling dataset [31] in Table 3 (baseline results and evaluation protocol taken from CViT [108]). UPT outperforms all compared methods, which are specifically designed for regularly gridded data.
Table 3: Comparison on the regular gridded small-scale ShallowWater-2D climate modeling dataset. UPT outperforms models that are specifically designed for regular grid data.
Model # Params Rel. L2 Error
DilResNet [93] 4.2M 13.20%
U-Net [86, 31] 148M 5.68%
FNO [51] 268M 3.97%
CViT-S [108] 13M 4.47%
UPT-S 13M 3.96%
D.4 ShapeNet-Car